# Single-Source Regular Path Querying in Terms of Linear Algebra

### v1

@@ 3 $\{G^{a}\}\leftarrow$ Boolean decomposition of the $\mathcal{G}$ adjacency matrix;
Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA represents query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$. Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be the input graph and $v_{s}\in V$ is a starting node. Then the algorithm 1 builds the following automata intersection:
@@
Core of the algorithm is a line 8 that performs one step of traversing two automata simultaneously. Result of the algorithm is a matrix $P_{|Q|\times|V|}$ such that
@@
An example of such a step is presented in figure 1. For graph and automaton with labels $\{a,b\}$, we visualise adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation. Solid-colored edges should be traversed simultaneously at the step. Dashed edges used to represent relation $M$: the edge connects two vertex that are simultaneously reachable. Initially, such an edge connects the initial state of the automaton and the starting vertex in the graph. After the step (one iteration of the main cycle), there are two such edges. One of them connects the final state of the automaton with vertex 5 of the graph. That means vertex 5 is reachable from the start vertex 3 by the path that farms a word acceptable by the automaton.
Hence, the set of reachable nodes can be easily obtained from the matrix $P$. Let $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$ be an evaluated 2-RPQ. Then by the main property of the algorithm 1 resulting matrix $P$ is such that
$v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P_{q_{F}v}=1\text{ for some }q_{F}\in Q_{F}.$
Procedure 2 computes the explicit representation of the query result. Namely, $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$. Formally, the following theorem holds.
@@
The proposed algorithm, represented in 1, computes matrix $P$ such that respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
@@
input : Matrix $P_{|Q|\times|V|}$, set of final states $Q_{F}$
output : Matrix $P^{F}_{1\times|V|}$
1
[…]
3 for $q\in Q_{F}$ do
       $F_{1q_{F}}\leftarrow 1$;
        // Setting up a matrix reducing $P$
4
5 end for
return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
Algorithm 2 Exporting reachable vertices subprogram
Proof of this theorem is straightforward induction by the paths length. Details are provided in appendix A.
Particularly, from the theorem, we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$. As far as $q_{F}$ is a final state of the NFA, $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\subseteq[\![\mathcal{N}]\!]$. Thus, $v$ reachable from $v_{s}$ by path satisfies the given regular constraint.
@@
The proposed single-source 2-RPQ algorithm can be used to solve the single-destination problem. Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in single destination query. Denote:
@@
The automaton $\mathcal{N}^{-}$ allows traversing paths in the opposite direction. It holds $\mathcal{R}^{-}=[\![\mathcal{N}^{-}]\!]$. Evaluating single-source reachability 2-RPQ on the reversed 2-RPQ $\mathcal{Q}^{-}$ gives using the 2-NFA $\mathcal{N}^{-}$:
@@
For performance reasons, it’s necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer non-zero entries. Associativity of the Boolean matrix multiplication can be used to perform less calculation. This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1:
$((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
In general, the first option is preferable due to the fact that 2-NFA for the query can be converted into the corresponding minimal deterministic finite automaton having no more than $|Q|$ non-zero cells in each $\mathcal{N}^{a}$ adjacency matrix due to the automaton determinism. This strictly leads to less expensive computation of the product with large $|V|\times|V|$ matrix.
@@ 5. Evaluation
The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) inside the LAGraph (Szárnyas et al., 2021) infrastructure [^6] The source code of the algorithm is available in our GitHub LAGraph fork: https://github.com/SparseLinearAlgebra/LAGraph/tree/rpq. Both sparse matrices and their transpositions are loaded in-memory due to the fact we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
Whereas we’re focusing on an efficient way of evaluating RPQs in-memory, there aren’t many candidates to compare with. Popular database management systems primarily focus on general availability and ensure availability of the concurrent access at the same time the suggested algorithm implementation is suited only for solving the reachability problem. We’ve selected the following systems as demonstrating the best performance and being the most related to our use case.
MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system supporting a synthetic way to carry out calculations without using the disk storage by caching the query data. It demonstrates state-of-the-art performance on evaluating regular path queries.
FalkorDB (ex. RedisGraph (Cailliau et al., 2019)) is an in-memory property-graph database also employing SuiteSparse:GraphBLAS for query evaluation. However, it uses OpenCypher modification and supports only a subset of regular path queries. To deal with this, we’ve carried out the measurements of cypher-compatible and non-cypher-compatible queries separately (e.g., it is impossible to evaluate repeated-path queries in the form of $(ab)^{*}$ in FalkorDB).
RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the in-memory linear-algebra based RPQ evaluation algorithm. The original work introduces a few variations of the adjacency matrix representation: $k^{2}$-trees offering less memory consumption and CSR/CSC formats providing better performance. We’ve chosen the last one to compare with since we’re aiming to compare the performance.
@@ 5.1. Dataset Description
For algorithm evaluation, we choose the Wikidata dataset from the snapshot provided in terms of MillenniumDB path query challenge (Farias et al., 2023). The resource contains both the graph and the set of 660 different 2-RPQs in SPARQL format taken from the Wikidata query log.
For systems that do not support RDF format, the dataset was converted to an edge-labelled graph. The processed graph has 610 402 396610402396610\,402\,396610 402 396 edges, 91 609 2549160925491\,609\,25491 609 254 vertices and 1 39513951\,3951 395 distinct labels.
SPARQL path queries has been converted to the corresponding minimized DFAs. For evaluating on FalkorDB the supported subset of regular path queries have been converted into Cypher queries of form MATCH ... COUNT (DISTINCT ...). The original dataset contains queries from all to all vertices and queries involving non-existing entities. They’ve been filtered out, resulting in 578 queries remaining.
5.2. Evaluation Results
@@
LA 2-RPQ MilenniumDB RPQ-Matrix FalkorDB
Mean C (ms) 83 327 348 1 625
Median C (ms) 2.5 4.4 4.2 18.8
Mean speedup C 1 3.9 4.2 19.6
Median speedup C 1 1.8 1.7 7.5
Mean NC (ms) 374 1 345 7 108
Median NC (ms) 37.4 69.8 269.6
Mean speedup NC 1 3.6 19.0
Median speedup NC 1 1.9 7.2
Memory (Gb) 18.1 6.9 50.0
BPT 31.9 12.1 85.8
Timeouts 0 1 15 53
@@
Table 1. Query execution results. Results are provided for cypher-compatible (C) and non-cypher-compatible (NC) queries separately. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
@@
We measure evaluation time per query with a timeout of 1 minute. All of the data required for query evaluation is pre-loaded into the memory. Queries are executed one by one.
We run MillenniumDB [^7] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 in a SPARQL mode with the maximum amount of CPU cores available. The query dataset is executed twice. The first run isn’t measured since it makes the database load the data into the memory. The query execution time is calculated as the internal optimizer plus execution time excluding the parser.
We execute FalkorDB v4.2.0888FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 querying through the python wrapper. The database is started without CPU core limitations and without query cache. We build an index over the vertices to be able to select the starting or destination vertex using it. The query execution time is based on the internal timer.
The RPQ-matrix algorithm is executed in CSR/CSC mode without core and memory consumption limitations. The consumed time is measured with the internal timer. The solution has been built out of the project repository [^9] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606.
Results of the measurements, both evaluation time and relative speedup, are provided in figure 2. Only queries expressible in Cypher are presented. The purple lines represent mean values, and the green lines represent medians. Note that relative speedup is represented as pointwise speedup for each query. So, mean and median calculated for per-query speedup. Numeric values of the mean and the median, the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are available within the table 1. Queries that have failed on any of the systems with timeouts are excluded from the statistics. The number of time outed queries shown in table 1.
We can see that while in median, speeding up for our solution is between 1.7 and 7.5, in mean we improve performance more than 19 times (relative to the most closest solutions RPQ-Matrix and FalkorDB). That means the proposed solution significantly outperforms other linear algebra based solutions on some queries (up to ×$\times 10^{5}$) while demonstrating comparable (and better in average) performance in the most cases. On the other hand, RPQ-Matrix demonstrates better performance on big number of queries. It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based solutions (proposed and RPQ-Matrix) to take the best from both of them.
Note that MillenniumDB is pretty stable: there is neither outstanding wins nor losses in performance. At the same time, the proposed solution is slightly better in average even though MillenniumDB caches query evaluation data, but our solution does not.
Memory consumption is presented on table 1. We don’t specify MillenniumDB memory consumption due to the fact it can process query in-memory only by caching data and doesn’t provide ways to preload whole dataset from the disk storage. RPQ-Matrix demonstrates the lowest memory usage. Memory consumption of the proposed solution is comparable with competitors.
We collected the top 10 queries with the worst and best speedup relative to all competitors. Collected data presented in tables 345. We can see that the weakness of our solution are relatively simple queries with a relatively small number of reachable vertices, like Q P131* ?x1 with less than 100 reachable vertices, or ?x1 (P138 / P106) Q with less than a million reachable vertices, where Q is a fixed vertex, and ?x1 is a variable. The forte of the proposed solution is a complex query (in terms of regexp complexity) with a huge number of reachable vertices.

### v2

@@ 8       $P\leftarrow P\operatorname{\operatorname{\oplus}}M$
Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be the input 2-NFA that represents the query and specifies the RL $\mathcal{R}=[\![\mathcal{N}]\!]$. Let $\mathcal{G}=\langle V,E,L,\lambda_{\mathcal{G}}\rangle$ be the input graph and $v_{s}\in V$ is a starting node. Then the algorithm 1 builds the following automata intersection:
@@
Result of the algorithm is a vector $P^{F}_{1\times|V|}$ for which $v\in[\![\mathcal{Q}]\!]_{SSR}\Leftrightarrow P^{F}_{1v}=1$ for some 2-RPQ $\mathcal{Q}=\langle\mathcal{G},\mathcal{R},v_{s}\rangle$.
The core of the algorithm is line 8 that performs one step of traversing two automata simultaneously and builds matrix $P_{|Q|\times|V|}$ such that:
@@
An example of such a step is presented in figure 1. For graph and automaton with labels $\{a,b\}$, we visualize adjacency matrices for both of these symbols, the matrix of relation $M$, and the process of new $M$ computation. The visualization uses solid-colored edges to represent transitions traversed simultaneously at the step, while dashed edges show the reachability relation $M$ — connecting vertices that are simultaneously reachable. Initially, $M$ contains just one edge linking the automaton’s initial state to the starting vertex in the graph. After one iteration of the main cycle, $M$ contains two edges. One of them connects the final state of the automaton with vertex 5 of the graph. This indicates that vertex 5 is reachable from the starting vertex 3 by the path that forms a word acceptable by the automaton.
The core idea of the algorithm can be summarized in a theorem that is proved using straightforward induction by the length of the paths. Details are provided in appendix A.
@@
The proposed algorithm, represented in 1, computes the matrix $P$ such that the respective relation $\mathcal{P}\subseteq Q\times V$ has the following property.
@@
In particular, from the theorem we can immediately conclude that for the given automaton $\mathcal{N}$, graph $\mathcal{G}$ and starting vertex $v_{s}$, $P_{q_{F}v}=1$ for $q_{F}\in Q_{F}$ iff exist two paths $\pi_{\mathcal{G}}$ to $v$ and $\pi_{\mathcal{N}}$ to $q_{F}$ such that $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}$. As far as $q_{F}$ is a final state of the NFA, $\omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\subseteq[\![\mathcal{N}]\!]$. Thus, $v$ reachable from $v_{s}$ by path satisfies the given regular constraint.
@@
The proposed single-source 2-RPQ algorithm can be used to solve the single-destination problem. Let $\mathcal{N}=\langle Q,\Sigma,\Delta,\lambda_{\mathcal{N}},Q_{S},Q_{F}\rangle$ be an NFA specifies the constraint in the single destination query. Denote:
@@
The automaton $\mathcal{N}^{-}$ allows one to traverse paths in the opposite direction. It holds $\mathcal{R}^{-}=[\![\mathcal{N}^{-}]\!]$. Evaluating single-source reachability 2-RPQ on the reversed 2-RPQ $\mathcal{Q}^{-}$ gives using the 2-NFA $\mathcal{N}^{-}$:
@@
For performance reasons, it is necessary to take into account the fact that sparse matrix multiplication works faster when matrices have fewer nonzero entries. The associativity of the Boolean matrix multiplication can be used to perform less calculations. This leads to the idea of two possible ways to calculate the product on line 8 in the algorithm 1: $((N^{a})^{T}\operatorname{\operatorname{\otimes}}M)\operatorname{\operatorname{\otimes}}G^{a}=(N^{a})^{T}\operatorname{\operatorname{\otimes}}(M\operatorname{\operatorname{\otimes}}G^{a}).$
In general, the first option is preferable due to the fact that 2-NFA for the query can be converted into the corresponding minimal deterministic finite automaton having no more than $|Q|$ non-zero cells in each $\mathcal{N}^{a}$ adjacency matrix due to the automaton determinism. This strictly leads to a less expensive computation of the product with the large $|V|\times|V|$ matrix.
When it comes to real-world implementation, it is important to note that it is also possible to implement BFS-based 2-RPQs without using a traversal matrix $M$ as described in algorithm 2. The complete description of such algorithm is available in appendix . For some sparse linear algebra libraries, it can be more efficient to deal with fewer distinct matrices than to have fewer nonzero entries, especially when there are very few of them. Such properties are common for executing simple regular path queries.
$(q,v)\in\mathcal{P}_{n}\Leftrightarrow\begin{cases}\exists\text{ 2-path }\pi_{\mathcal{G}}\text{ of length }\leq n\text{ in }\mathcal{G}\text{ from }v_{s}\text{ to }v\\ \exists\text{ path }\pi_{\mathcal{N}}\text{ of length }\leq n\text{ in }\mathcal{N}\text{ from }q_{s}\in Q_{S}\text{ to }q\\ \omega^{\leftrightarrow}_{\mathcal{G}}(\pi_{\mathcal{G}})\cap\omega_{\mathcal{N}}(\pi_{\mathcal{N}})\neq\operatorname{\varnothing}.\end{cases}$
output : Vector $P^{F}$ of size $1\times|V|$
1
4 $P_{|Q|\times|V|}\leftarrow P$ s.t. $P_{qv}=1$ if $q\in Q_{S}$, $v=v_{s}$, otherwise 00;
5 $F_{1\times|Q|}\leftarrow F$ s.t. $F_{1}q=1$ if $q\in Q_{F}$, otherwise 00;
6
7while $P$ changes do
       $P\leftarrow P\oplus\bigoplus_{a\in\Sigma^{\leftrightarrow}\cap L^{\leftrightarrow}}((N^{a})^{T}\operatorname{\operatorname{\otimes}}P\operatorname{\operatorname{\otimes}}G^{a})$;
        // Update $\mathcal{P}$
8
9 end while
10
return $P^{F}=F\operatorname{\operatorname{\otimes}}P$
Algorithm 2 Single Source 2-RPQ using linear algebra without an extra traversal matrix
Although it may seem that this algorithm should perform some extra calculations because $P$ contains more entries than $M$ on each step within algorithm 1, this way of implementing the proposed algorithm can be handy. Some sparse linear algebra libraries perform better when handling fewer distinct matrices, even if that increases the total number of non-zero entries. Such properties are common for executing simple regular path queries.
4.1. Comparison With Diego Arroyuelo Algorithm
Diego Arroyuelo et al. in (Arroyuelo et al., 2023a) propose another linear algebra-based 2-RPQ evaluation algorithm called RPQ-matrix. It directly translates 2-RE to Boolean matrix operations rather than evaluating step-by-step traversal over the graph and the automaton with capability to employ single-source and single-destination 2-RPQs. The algorithm consists of the following steps.
  1. (1)
    Build an abstract syntax tree (evaluation plan) representing the desired regular expression in which leaves represent labels and other nodes represent operations such as concatenation, Kleene star, and conjunction.
  2. (2)
    Match nodes with matrices: each leaf label is matched with an adjacency matrix representing it, and every inner node is matched with a matrix that can be computed based on the child matrices depending on the operation the node represents.
  3. (3)
    Optimize the evaluation plan by using the provided source or destination vertex.
  4. (4)
    Compute the matrix representing the root element with operations reordering applied to optimize computations.
We conduct experiments on several graph database management systems and compare them to RPQ-matrix and to the proposed algorithm in order to investigate the efficiency of different linear algebra-based approaches.
@@ 5. Evaluation
The proposed single-source 2-RPQ algorithm is implemented using the SuiteSparse:GraphBLAS library (Davis, 2019b) within the LAGraph (Szárnyas et al., 2021) infrastructure [^5] Regular Path Query algorithm in LAGraph repository: https://github.com/GraphBLAS/LAGraph/blob/stable/experimental/algorithm/LAGraph_RegularPathQuery.c. Both sparse matrices and their transpositions are loaded in memory due to the fact that we need both representations to efficiently evaluate 2-RPQ and traverse the graph in both directions.
Whereas we are focusing on an efficient way of evaluating RPQs in memory, there are not many candidates to compare. Popular database management systems primarily focus on general availability and ensure availability of the concurrent access at the same time the suggested algorithm implementation is suited only for solving the reachability problem. We have selected the following systems as the most effective and the most related to our use case.
RPQ-matrix (Arroyuelo et al., 2023a) is another implementation of the linear algebra-based RPQ evaluation algorithm. The original work introduces a few variations of the adjacency matrix representation: $k^{2}$-trees offering less memory consumption and CSR/CSC formats providing better performance. We have chosen the last one to compare since we are aiming to compare the performance.
RPQ-matrix (GrB) [^6] SuiteSparse:GraphBLAS-based implementation of RPQ-matrix: https://github.com/suvorovrain/rpq-matrix/tree/gbmod is an adapted version of Diego Arroyuelo et al. RPQ-matrix algorithm where matrix representations and operations are substituted with SuiteSparse:GraphBLAS equivalents in order to analyse performance impact of basic primitives implementation.
MillenniumDB (Vrgoc et al., 2021) is a graph-oriented database management system with RDF-model and SPARQL support. It supports a synthetic way to carry out calculations without using disk storage by caching query data. MillenniumDB demonstrates state-of-the-art performance on evaluating regular path queries.
FalkorDB (previously RedisGraph (Cailliau et al., 2019)) is an in-memory property graph database that also employs SuiteSparse:GraphBLAS for query evaluation. However, it uses OpenCypher modification and supports only a subset of regular path queries (e.g., it is impossible to evaluate repeated-path queries in the form of $(a\;b)^{*}$). To deal with this, we have carried out the measurements of cypher-compatible and non-cypher-compatible queries separately.
Blazegraph (Vrgoc et al., 2021) is another graph-oriented database management system using RDF data model and SPARQL query language. It is used by the Wikidata project and is based on more classical B-trees.
@@
5.1. Implementation Details
As mentioned above, it is important for better performance to take into account that there are two different possible implementations of the BFS-based 2-RPQ algorithm: one involving less different matrices and one with less dense matrices. For SuiteSparse:GraphBLAS, the first approach turns out to be faster if the number of entries in matrices is small, and the latter is better for denser matrices.
The most straightforward way to combine these two approaches to achieve better performance is to switch from the first approach to the second one during the traversal if the resulting matrix starts having some constant amount of non-zero entries. This constant can be empirically determined for the graph [^7] For the studied datasets the most suitable value is $100$.. This allows the BFS-based algorithm to provide strong performance on simple queries with very few answers and on analytical queries involving a lot of resulting vertices at the same time.
@@ 5.2. Dataset Description
To compare performance of different approaches we start from evaluating benchmarks on large real-world datasets. For algorithm evaluation, we choose the Wikidata dataset from the snapshot provided in terms of MillenniumDB path query challenge (Farias et al., 2023). The resource contains both the graph and the set of 660 different 2-RPQs in SPARQL format taken from the Wikidata query log. The second dataset is Yago-2S evaluated with 7 different complex queries taken from (Abul-Basher et al., 2017).
We also want to compare different linear algebra-based approaches and determine the best of them for different query kinds. Real-world datasets are not suitable for it due to complex graph topology. Instead, we employ the synthetic RPQBench dataset generator (Wang et al., 2025) for controlled performance evaluation across query categories. The structure of the graph ensures reproducible and predictable results when similar queries are executed. The original generator produces arbitrary sized RDF datasets and offers $10$ different query kinds without starting or final nodes specified in SPARQL format. We generate a graph and supply these query kinds with randomly generated source and destination vertices.
For systems that do not support the RDF format, the datasets have been converted to an edge-labeled graph. Original SPARQL path queries have been deprefixed, converted to the corresponding minimized DFAs. Queries have been converted to Cypher queries of the form MATCH ... COUNT (DISTINCT ...) when it is possible. Queries without starting or final vertex, broken queries involving missing entities have been removed. Final dataset statistics can be summarised as follows.
Wikidata
  • $610$ million edges, $91$ million vertices, $1400$ distinct labels.
  • $578$ queries filtered out from $660$ from the query dump.
Yago-2S
  • $610$ million edges, $91$ million vertices, $1400$ distinct labels.
  • $7$ complex queries taken from (Abul-Basher et al., 2017).
RPQBench
  • Synthetic dataset, $610$ million edges, $91$ million vertices, $1400$ distinct labels.
  • $20$ different query kinds supplied $1000$ randomly generated source/destination vertices.
  • Trivial queries are filtered out (e.g. $a^{*}$ has $\geq 1$ answer).
5.3. Evaluation Scenario
((a)) Queries which are supported and succeeded on all of the competitors
((b)) Queries which are not supported in Cypher or timed out on FalkorDB or Blazegraph
@@ 5.3. Evaluation Scenario
Table 1. Wikidata query execution results. The results are supplied for simple and for complex (C) queries separately. For in-memory algorithms the memory consumption for whole dataset and bytes-per-triple values (BPT) are presented. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
LARPQ RPQ-matrix RPQ-matrix (GrB) MillenniumDB FalkorDB Blazegraph
Total, ms 24 403 166 882 34 891 276 527 460 606 403 738
Mean, ms 49.8 340.6 71.2 564.3 940.0 824.0
Median, ms 2.0 3.3 1.1 13.7 17.6 128.0
Mean speedup 1.00 0.15 0.70 0.09 0.05 0.06
Median speedup 1.00 0.60 1.74 0.15 0.11 0.02
Total C, ms 35 061 508 378 288 766 639 193
Mean C, ms 427.6 6 199.7 3 521.5 7 795.0
Median C, ms 169.1 372.9 142.9 3 105.6
Mean speedup C 1.00 0.07 0.12 0.05
Median speedup C 1.00 0.45 1.18 0.05
Memory, Gb 9.2 6.9 9.2
BPT 16.3 12.1 16.3
Table 2. Yago-2S query execution results. Mean and median speedup relatively the proposed solution is a ratio of mean and median over query set for respective systems
LARPQ RPQ-matrix
RPQ-matrix
MillenniumDB Blazegraph
Total, ms 377 174 202 7 025 24883
Mean, ms 54 25 29 1 004 3555
Median, ms 69 25 38 985 3808
Mean speedup 1.00 2.17 1.87 0.05 0.02
Median speed up 1.00 2.71 1.81 0.07 0.02
Memory, Gb 0.5 0.4 0.5
BPT 11.3 9.8 11.3
@@
We measure query evaluation time with a 1-minute timeout, preloading all required data into memory, so data preprocessing time is not included. Queries execute sequentially in isolation.
RPQ-Matrix [^8] RPQ-Matrix repository we use for the evaluation: https://github.com/adriangbrandon/rpq-matrix/tree/34fc2240a7c8069f7d6a39f1c75176edac4fe606 configuration.
  • CSR/CSC mode as the most performant one.
  • No CPU/cores limitations.
  • No memory limitations.
  • Timing: internal timer.
RPQ-matrix (GrB) [^9] RPQ-matrix implemented using SuiteSparse:GraphBLAS repository: https://github.com/suvorovrain/rpq-matrix/tree/gbmod configuration.
  • Pre-caching CSR/CSC matrices.
  • No CPU/cores limitations.
  • No memory limitations.
  • Timing: internal timer.
MillenniumDB [^10] MillenniumDB repository we use for the evaluation: https://github.com/MillenniumDB/MillenniumDB/tree/5190c0d9b07ca681328495b69c715af792513775 configuration.
  • SPARQL mode.
  • Maximum CPU cores.
  • Two execution runs (first warm-up excluded).
  • Timing: internal timer for query optimization + execution (parser excluded).
FalkorDB (v4.2.0) [^11] FalkorDB v4.2.0: https://github.com/FalkorDB/FalkorDB/tree/v4.2.0 configuration.
  • Accessed via Python wrapper.
  • No CPU/cores limitations.
  • No query caching.
  • Vertices index for efficient start/end vertex selection.
  • Timing: internal database timer.
Blazegraph (2.1.5) [^12] Blazegraph v2.1.5: https://github.com/blazegraph/database/tree/BLAZEGRAPH_RELEASE_2_1_5 configuration.
  • Accessed via Python wrapper.
  • No CPU/cores limitations.
  • No memory limitations.
  • Timing: external timer.
  • Queries are executed one by one with extra heatup to avoid out-of-memory.
All benchmarking scripts, configurations, and instructions are available in GitHub repository [^13] Benchmark evaluation repository: https://github.com/SparseLinearAlgebra/la-rpq.
5.4. Real-World Graph Querying Results
Resulting per-query evaluation time is provided in Figure 2(a) for queries that are supported on all of the competitors. Remaining complex queries ended up with a timeout on slower competitors or not expressible in Cypher are represented separately in Figure 2(b). The dotted lines represent means and the straight lines represent medians. These numeric values of and total execution time are available in the table 1 for the relatively simple and complex queries (C) separately. For linear algebra-based competitors the size of the dataset loaded in-memory, and memory byte-per-triple (BPT) memory consumption values are provided.
For simple queries, LARPQ demonstrates the best mean with a speedup of 1.7×1.7$\times$1.7 × to 18.9×18.9$\times$18.9 ×. However, RPQ-matrix implemented with SuiteSparse:GraphBLAS demonstrates the best median time that is better 1.7×1.7$\times$1.7 × than that of LARPQ. It means that RPQ-matrix evaluates some of the queries faster whereas the BFS-based algorithm works better in general cases. Our hypothesis is that the proposed algorithm does not utilise the sparsity of some edge kinds enough. Both linear algebra-based approaches demonstrate better time in mean an in median in comparison to all of database management systems.
For complex queries, relations between different competitors are the same except that LARPQ mean is drastically lower than other competitors being 14.5×14.5$\times$14.5 ×, 8.3×8.3$\times$8.3 ×, 18.3×18.3$\times$18.3 × less than those of the competitors. It means the proposed algorithm is capable of executing the most complex out of the complex algorithm faster than other approaches.
The results for the Yago-2S dataset are presented in table 2. FalkorDB is excluded since every of the $7$ queries have resulted with a timeout.
Both RPQ-matrix algorithms demonstrate better time than the proposed LARPQ algorithm. It is likely due to the structure of the query. All of them use $a\;b\;c^{+}\;d^{+}$ patterns. Iterative structure of the BFS-based approach does not utilize efficient evaluation order that is important for such long queries of simple operations.
It is an interesting question for future research whether it is possible to combine ideas from two linear algebra based-solutions (proposed and RPQ-Matrix) to take the best from both of them.
5.5. Synthetic Graph Querying Results
We perform a detailed comparison of competitors for executing different query kinds by running a synthetic RPQBench. Its evaluation results are available in the table 3. Each row represents distinct query kind. Total execution time of the randomly generated queries are presented in seconds for each competitor separately. Due to the lower overall performance demonstrated on the real-world datasets, MillenniumDB, FalkorDB, and Blazegraph are excluded from subsequent comparisons. We also provide edge statistics for each label kind in table 4.
Table 3. RPQBench dataset evaluation time of queries with randomly generated sources and destinations per each query kind in seconds
Query pattern, single-source
(S) or single-destination (D)
LARPQ RPQ-matrix
RPQ-matrix
1 $a\;b\;c$, (S) 51 83 11
2 $a\;b\;c$, (D) 32 11 12
3 $(a\;b\;c)\mid(c\;d\;d)$, (S) 59 89 24
4 $(a\;b\;c)\mid(c\;d\;d)$, (D) 47 14 22
5 $d^{*}$, (S) 22 21 37
6 $d^{*}$, (D) 19 19 23
7 $d^{*}\;e$, (S) 21 12 18
8 $d^{*}\;e$, (D) 6 3 5
9 $d^{+}$, (S) 23 14 29
10 $d^{+}$, (D) 18 14 20
11 $(a\;b)^{*}$, (S) 1 350 30 156 4 904
12 $(a\;b)^{*}$, (D) 3 619 15 706 2 978
13 $f\;g\;(d\mid e)$, (S) 92 700 5 917 7 745
14 $f\;g\;(d\mid e)$, (D) 2 193 301 486
15 $f\;g\;(d\mid e)^{*}$, (S) 149 644 2 870 709 2 053 090
16 $f\;g\;(d\mid e)^{*}$, (D) 6 167 810 1 053
17 $(c\mid g)\;(d\mid e)$, (S) 36 3 7
18 $(c\mid g)\;(d\mid e)$, (D) 16 173 017 41 713
19 $(c\mid g)\;(d\mid e)^{*}$, (S) 930 55 232
20 $(c\mid g)\;(d\mid e)^{*}$, (D) 99 955 891 229 035
Table 4. RPQBench edge stats per label
Edge label Count Edge label Count
$a$ 343 660 $e$ 36
$b$ 4 209 447 $f$ 4 928 456
$c$ 114 742 222 $g$ 223 656
$d$ 186
As it is observed, the execution time of the proposed BFS-based algorithm for simple queries such as $d^{*}$, $d^{+}$, and $d^{*}e$ is quite similar to that of other linear algebra-based implementations. RPQ-matrix ×$\times 2$ speedups are likely to happen due to the CSR/CSC matrix implementation since RPQ-matrix (GrB) demonstrates evaluation time close to LARPQ.
[…]
For the remaining queries, such as 1–4, the proposed algorithm demonstrates a relative slowdown of 2×2$\times$2 × to 4×4$\times$4 ×, since it does not utilize an efficient evaluation order. For queries 14, 16, 17, and 19, LARPQ is 4×4$\times$4 × to 17×17$\times$17 × slower than the competitors because it does not exploit the fact that edges with labels $d$ or $e$ are very rare.
Conclusion about various linear algebra-based algorithms might be summarized as follows.
  • LARPQ is a better choice for queries that involve compound operations over labels corresponding to many edges and not having source/destination vertices close to them.
  • LARPQ algorithm tends to be more stable whereas RPQ-matrix might demonstrate drastic slowdown for some kinds of complex queries.
  • Both algorithms are efficient enough for simple queries.
  • RPQ-matrix is a better choice if the query contains rare labels, long concatenations or disjunctions allowing to perform optimizations.